Hendonism was pandemic as .87 of the students were hedonists. If 1914 students were hedonists how many students were there in all?
Answer by rajagopalan(174)

0.87 of the students were hedonists=1914
87 % of students =1914
then 1% of students=1914/87=22
So Total Students =100%= 22*100=2200